OpenStudy (anonymous):

A drawer contains n = 5 different and distinguishable pairs of socks (a total of ten socks). If a person (perhaps in the dark) randomly selects four socks, what is the probability that there is no matching pair in the sample total combinations 10C4= 210 Of the 5 matching pairs we have 28 * 5 ways that we can match only one pair when four are selected (blue, blue, x,y) 8C2= 28 (red,red,x,y) . . . so we get 5*28 combinations where socks match so combinations with no match 210 - 140= 70 probabiliy no match = 70/210 is this righ??

OpenStudy (anonymous):

one third?

OpenStudy (anonymous):

i get a different answer, but i could be wrong

OpenStudy (anonymous):

pick a sock it is some color pick another sock the probability it is different color is \(\frac{8}{9}\) the probability that the next is different from the first two is \(\frac{6}{8}\) and the probability that the fourth is different from the first three is \(\frac{4}{7}\)

OpenStudy (anonymous):

take the product for your answer, it is not one third i don't think

OpenStudy (reemii):

I agree with satellite73, with (8*6*4)/(9*8*7) as anwer.

OpenStudy (anonymous):

for sure there are 140 , where socks match right? so now we would have to think if there is anything else missing I guess

OpenStudy (anonymous):

it does make sense

OpenStudy (reemii):

you forgot the cases: (blue, blue, red, red) , ...

OpenStudy (anonymous):

couldn't they be in the (blue,blue,x,y) x and y could be anything including other matching socks no?

OpenStudy (reemii):

you wrote 8C2. this does .. oops you're right.

OpenStudy (reemii):

but there's a problem linked to this. (blue, blue, x,y) with (x,y)=(red,red) means that you count (blue, blue, red,red) here once, but you will count it again when you look at (red, red, x,y) and (x,y)=(blue,blue). Therefore, you count too many stuff in the matching probability, therefore the probability of "no match" is too small.

OpenStudy (reemii):

@arykk

OpenStudy (anonymous):

true you are right once again. so there fore we would have to take away those . in that case 210 - 140-10 = 60 combinations where socks don't match. so the probability would be 60/210 =2/7 = .2857 this would make sense how come it isn't the same as satellite73 approach? (8*6*4)/(9*8*7) = .3809

OpenStudy (anonymous):

@satellite73 @reemii

OpenStudy (reemii):

it's rather 210-140 + something

OpenStudy (reemii):

because it's 210 - (140 - something) (140 was too big)

OpenStudy (anonymous):

this approach is too confusing for me

OpenStudy (reemii):

number of (x,x,y,y) is 5C2 = 10. as you said. it's ok: (210-140+10)/210 = satellite's answer = 0.38095... !

OpenStudy (reemii):

@arykk I think you can assume you solved the problem your way too.

OpenStudy (anonymous):

@reemii thanks!!